Circular No. 8343

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 K1 (CATALINA) An apparently asteroidal object reported by the Catalina Sky Survey

(discovery observation given below), and posted on the NEO Confirmation Page, has been found by T. Spahr to show a weak coma with a possible very faint extension to the west on CCD images obtained by J. B. Battat and himself (Mt. Hopkins 1.2-m reflector) on May 24.4 UT; the FWHM of the comet image is 20 percent larger than other stars of similar brightness in the field. R. Stoss reports that CCD images taken by S. Sanchez, J. Nomen, and himself on May 24.1 (0.30-m reflector, Mallorca) shows the object to have a softer image than surrounding field stars.

2004 UT  α_{2000}  δ_{2000}  Mag.
May 21.40758  20^{h}43^{m}06.^{s}88  −4^{o}17′23′′.0  18.6

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-K50.

T = 2005 July 4.680 TT  ω = 97.^{o}280
Ω = 326.888 }2000.0
q = 3.41992 AU  i = 153.701

SUPERNOVA 2004bu IN UGC 10089 Further to IAUC 8341, T. Boles reports his discovery, on unfiltered

CCD images taken on May 18.981 and 19.896 UT, of an apparent supernova (mag 17.1) located at α = 15^{h}54^{m}52^{s}.03, δ = +21^{o}07′11′′.0 (equinox 2000.0), which is ≈ 1′′.2 west and 10′′.5 north of the center of UGC 10089. SN 2004bu is not present on Boles’ images from Mar. 20 and Jan. 29 (limiting mag 19.5) or on Digitized Sky Survey red (1992) or blue (1993) plates.

SUPERNOVA 2004br IN NGC 4493 C. L. Gerardy, University of Texas at Austin, reports that a low-

resolution optical spectrogram (resolution 300; range 500–1000 nm) of SN 2004br (cf. IAUC 8340), obtained on May 18.20 UT with the 9.2-m Hobby/ Eberly Telescope by B. Roman and F. Deglman, shows it to be an unusual type-Ia supernova, similar to the spectrum of SN 2000cx at three days before maximum light (Li et al. 2001, PASP 113, 1178). Adopting the NED redshift of NGC 4493 (6949 km/s), the expansion velocity of the (rather weak) Si II feature (rest wavelength 635.5 nm) is 9500 km/s.

2004 May 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green